### v7

@@ 3.3 Stage-2
In Stage-2, the proposed method is illustrated in Figure 5. Initially, the original input images are split into three RGB single-channel images and one grayscale image. Each of the single-channel images is input into an independent generator, resulting in the use of four generators in Stage-2 of the proposed method. These four generators train the model separately, enabling this method to remove background information from the localized image blocks while extracting colored foreground details. Additionally, all single-channel images share the same discriminator to distinguish between the generated image and its corresponding ground truth image.
@@ 3.3.1 Loss functions of GAN
To ensure a more robust convergence of the loss function, we follow the improved Wasserstein generative adversarial network with gradient penalty (WGAN-GP) [37] used by Suh et al. [31] applied to the objective function. Experiments by Bartusiak et al. [38] have demonstrated that the binary cross-entropy (BCE) loss yields superior performance compared to the $L1$ loss in binary classification tasks. Therefore, we chose the BCE loss instead of the $L1$ loss used in the Pix2Pix GAN [26]. The loss function of the WGAN-GP, which incorporates the BCE loss, is formulated as follows:
@@
where the penalty coefficient is $\alpha$, and the uniform sampling along a straight line between the ground truth distribution $P_{y}$ and the point pairs of the generated data distribution is $P_{\hat{y}}$. $\lambda$ is used to control the relative importance of different loss terms. In the discriminator, the generated image is distinguished from the real image by the loss function $\mathbb{L}_{D}$ in Eq. (5). In the generator, the distance between the generated image and the ground truth image in each color channel is minimized by the loss function $\mathbb{L}_{G}$ in Eq. (6).
3.3.2 Network architecture
Given the unpredictability of color document degradation, this work adopts four independent adversarial networks to extract text information from diverse colored backgrounds. This method mitigates color interference during document image binarization.
For the network proposed in Stage-2, four independent generators apply the same encoder-decoder architecture. The encoder extracts features, while a shared decoder is employed by multiple encoders of varying depths, and subsequently the features are concatenated. A skip connection is used between the encoder and the decoders. The encoder facilitates downsampling and context extraction, while the decoders undertake upsampling and amalgamate both upscaled and downsampled features. Specifically, this work employs EfficientNet-B6 [36] as the encoder, which is a lightweight CNN known for its performance in image classification tasks. For the discriminator, we choose the one [26, 29, 30, 39, 40, 41] used by Suh et al. [31] due to its superior generalization capabilities. Additionally, instead of U-Net [21] used by Suh et al. [31], this work uses its improved model U-Net++ [34, 35] for semantic segmentation, which can further improve the model performance.
The detailed procedure for training the proposed GAN models is illustrated in Algorithm 1. The training process is conducted in our training environment, which is described in Section 4.3, encompassing optimizer selection, learning rate configuration, and loss function definition. The algorithm demonstrates that the GAN model iteratively updates the loss function based on the authenticity assessment of the generated image by the discriminator. This iterative process enhances the ability of the generator to produce more real images.
@@ 3.4 Stage-3
In Stage-3, a multi-scale adversarial network is employed to generate images for both local and global binarization results, enhancing the distinction between the background and text with higher accuracy. Contrary to Stage-2, which predominantly relies on local prediction using the image patches ($224\times 224$), Stage-3 conducts global prediction on the resized input image ($512\times 512$). This method reduces the contextual information loss in text caused by the local predictions of Stage-2.
Figure 6 illustrates the process of two independent discriminators in Stage-3. Different from Stage-2 using 8-bit images for local prediction, Stage-3 uses 24-bit images as the input of the generator. Additionally, Figure 6 clarifies that the left-side input image comprises the image patches from the Stage-2 output, intended for local prediction. Conversely, the right-side input image represents the resized image from the original input for global prediction. In the binarization process of Stage-3, the network architecture of the generator remains consistent with Stage-2, differing only in the number of input channels.
@@ 4.1 Datasets
4.1.1 DIBCO
The Document Image Binarization Competition (DIBCO) provides several datasets: DIBCO 2009 [42], H-DIBCO 2010 [43], DIBCO 2011 [11], H-DIBCO 2012 [44], DIBCO 2013 [12], H-DIBCO 2014 [13], H-DIBCO 2016 [14], DIBCO 2017 [15], and H-DIBCO 2018 [16]. These datasets comprise both grayscale and color images, including machine-printed as well as handwritten content. Participants in the competition were tasked with designing algorithms capable of extracting binarized text images from both handwritten and machine-printed images within these datasets.
@@ 4.1.3 SMADI
The Synchromedia Multispectral Ancient Document Images Dataset (SMADI) [1] consists of 240 multispectral images of 30 authentic historical handwritten letters. These letters, dating from the 17th to the 20th centuries, were written using iron gall ink. The original documents are deposited at the BAnQ Bibliotheque et Archives nationales du Quebec. A total of 240 images of real documents were meticulously captured, calibrated, and registered within this dataset. Each of these images was imaged with a CROMA CX MSI camera, resulting in 8 images per document.
@@ 4.1.4 Bickley Diary Dataset
The Bickley Diary Dataset (BD) [46] has been generously donated to the Singapore Methodist Archives by Erin Bickley. This dataset comprises diaries in which the images have been affected by both light and fold damages, thereby complicating the process of identification.
@@ 4.2 Evaluation Metric
In comparison with other traditional and SOTA methods for document image binarization, based on the case that DIBCO datasets do not provide OCR output, we use the following five evaluation metrics:
@@ 4.2.1 F-measure (FM)
where the Recall is defined as $\frac{TP}{TP+FN}$; the Precision is defined as $\frac{TP}{TP+FP}$, and TP, FP, and FN represent the true positive, false positive, and false negative values, respectively.
@@ 4.2.2 Pseudo-F-measure (p-FM)
where the Pseudo-Recall (${\rm Recall_{ps}}$) and the Pseudo-Precision (${\rm Precision_{ps}}$) both use the distance weights relative to the ground truth image profile to generate a new weighted ground truth image.
@@ 4.2.3 Peak signal-to-noise ratio (PSNR)
where ${\rm V}$ represents the difference between the foreground and background, and the Mean Square Error (MSE) is defined as:
@@ 4.2.4 Distance reciprocal distortion (DRD)
${\rm DRD}$ represents a measure of visual distortion in the binary image, which is shown in the following equation:
@@
where ${\rm NUBN}$ corresponds to the number of uneven $8\times 8$ blocks in the ground truth image, and ${\rm DRD_{k}}$ is defined as:
@@
where ${\rm DRD_{k}}$ indicates the distortion of the ${\rm k_{th}}$ flipped pixel, calculated using a $5\times 5$ normalized weight matrix ${\rm N_{w}}$. ${\rm B_{k}}$ refers to the pixel of the binary image, and ${\rm G_{k}}$ denotes the pixel of the ground truth image.
@@ 4.2.5 Avg-Score (Avg)
The Peak Signal-to-Noise Ratio (PSNR) provides a straightforward and intuitive quantitative metric for comparing the generated image with the ground truth image. Although it can be used as a supplementary metric, it should not be the only evaluation measure. The F-measure and p-F-measure are used to assess the accuracy of different methods, but a higher value in these metrics does not always correspond to a higher PSNR value. In our experiments, it is common to observe that a method achieves the SOTA PSNR value, while the F-measure and p-F-measure values are lower than those of other models. To solve this problem, we followed the Avg-Score metric used by Jemni et al. [47] to compare the combined performance of different methods, as shown in the following equation:
@@
Table 1: Comparison of the PSNR values of the images from different training sets that have been processed using different methods with the corresponding ground truth images.
@@
Original Input 63.2879 61.3446 61.8542 57.6995 60.7355 61.2175 61.0232
DWT 64.1154 71.4422 64.8151 53.7037 51.5094 60.1374 60.9539
DWT&Norm 80.4286 86.1904 82.4625 74.2314 83.6403 78.5337 80.9145
Table 2: Quantitative comparison (FM/p-FM/PSNR/DRD/Avg) of the models trained with different patch sizes ($224\times 224$/$256\times 256$) on multiple benchmark datasets.
DIBCO 2011 DIBCO 2013 H-DIBCO 2014 H-DIBCO 2016 DIBCO 2017 DIBCO 2018 Mean Values
@@
Figure 7: Visualization examples of the results of the proposed method at each stage. (a) the original input images, (b) the LL subband images obtained by Haar wavelet transform and normalization (Stage-1), (c) the enhanced images produced using our image enhancement method (Stage-2), (d) the binarized images obtained by combining local and global predictions (Stage-3), (e) the ground truth images.
Table 3: Comparison of performance evaluation of the proposed method and other methods for document binarization on multiple benchmark datasets. Best and 2nd best performance are in red
and blue colors, respectively. The data for the experimental results of other methods are obtained from [31, 48].
@@
(f) DIBCO 2018
Figure 8: Example of document image binarization results from input image HW1 of DIBCO 2011 dataset by different methods: (a) original input, (b) Otsu [6], (c) Niblack [5], (d) Sauvola [7], (e) Vo [23], (f) He [24], (g) Suh [31], (h) Ours, (i) ground truth.
Figure 9: Example of document image binarization results from input image HW5 of DIBCO 2013 dataset by different methods: (a) original input, (b) Otsu [6], (c) Niblack [5], (d) Sauvola [7], (e) Vo [23], (f) He [24], (g) Suh [31], (h) Ours, (i) ground truth.
@@ 4.3.1 Data preparation
The performance of a neural network depends on the quality and quantity of its training data. In general, the more training data enhances the better the performance of the neural network. To compare the performance of the proposed method with other methods, we follow the strategy described in [23, 24, 31] for constructing the training and test sets. Specifically, the training set includes a total of 143 images from DIBCO 2009 [42], H-DIBCO 2010 [43], H-DIBCO 2012 [44], PHIBD [45], SMADI [1], and BD [46] dataset. The test sets comprise a total of 82 images from DIBCO 2011 [11], DIBCO 2013 [12], H-DIBCO 2014 [13], H-DIBCO 2016 [14], DIBCO 2017 [15], and H-DIBCO 2018 [16].
To ensure a fair comparison of performance among all methods, we used the same data augmentation for our method and other SOTA methods. During the training process of the local prediction network, this work splits the input images into patches with $224\times 224$ in size. These patches were sampled with scale factors of ${0.75,1.25,1.5}$, subsequently resized to $224\times 224$, and rotated by 270°. In contrast, for training the global prediction network, this work employed scaling without rotation for data augmentation, using a scaling size of $512\times 512$. To diversify the training data, rotation augmentation was applied to the global prediction, involving rotations of 90°, 180°, and 270°, as well as horizontal and vertical flips.
@@ 4.3.2 Training
The experiments were conducted using the PyTorch framework in Python. We use NVIDIA RTX 3090 GPUs for all the following experiments.
In this study, we employ the generator comprising EfficientNet-B6 [36] with UNet++ [34, 35]. Regarding the training hyperparameters, Stage-2 and Stage-3 share similar settings, differing primarily in the number of training epochs: 150 epochs for global prediction and 10 epochs for the remaining networks. The choice of optimizer was Adam [49], with a learning rate set at $2\times 10^{-4}$. Additionally the generator employed $\beta_{1}=0.5$, while the discriminator used $\beta_{2}=0.999$.
@@ 4.3.3 Pre-training
To enhance training efficacy due to limited data availability, we employed EfficientNet-B6 [36] as the encoder in the generator, utilizing weights pretrained on the ImageNet dataset [50].
@@ 4.4 Comparison with other Methods
In our previous conference paper [51], we compared various deep learning-based methods and found that some of these methods [10, 28, 29] employed a “leave-one-out” strategy for selecting the training set. For instance, when the DIBCO 2013 dataset is used as the test set, all other datasets are utilized as the training set. Although our method still demonstrates superior performance at the SOTA level in such case [51], we consider it unfair to compare them directly because our model uses a smaller training set. Therefore, this work excludes these methods from this paper and only retains those that use the exact same training set for comparison.
@@ 4.5 Image Enhancement Experiment
In Stage-1, we applied DWT and normalization to the original input images. While mathematical theory [52] suggests that these processed images can retain contour information and reduce noise more efficiently, it is important to demonstrate their impact on experimental results. The previous conference paper [51] on this work presented ablation experiments on DWT and normalization. Based on this, it is better to intuitively show that, compared to the original input images, the DWT and normalization processed images are closer to the ground truth images.
In Stage-1 of this work, we set the threshold to 0.3 to binarize different single-channel images to obtain corresponding binarized images. In the experiment of this section, we applied the above image binarization to the original input images, the images processed with DWT, and the images processed with both DWT and normalization. Subsequently, we calculated the PSNR values of the binarized images compared to the ground truth images using these three methods, and the results are presented in Table 2. The PSNR mean values of the original input images are 61.0232. In contrast, the PSNR mean values of the images after DWT processing are slightly lower at 60.9539. However, the PSNR mean values of the images processed with both DWT and normalization are 80.9145, which means they are closer to the ground truth images. This demonstrates that it is more reasonable to choose the DWT and normalization processed image rather than the original input image and the DWT processed image.

### v8

@@ 3.3 Stage-2
In Stage-2, the proposed method is illustrated in Figure 6. Each single-channel patch image is input into an independent generator, so four generators are used in Stage-2 of the proposed method, which enable the method to extract color foreground information while removing background noise from the patch images. In addition, all single-channel images share the same discriminator to distinguish between the generated images and the corresponding Ground-Truth images.
@@ 3.3.1 Loss functions of GAN
To ensure a more robust convergence of the loss, we follow an improved loss function of the Wasserstein generative adversarial network with gradient penalty (WGAN-GP) [39] used by Suh et al. [31], where WGAN-GP loss incorporates the BCE loss. Since Bartusiak et al.’s [40] experiments demonstrate that the binary cross-entropy (BCE) loss outperforms the $L1$ loss in binary classification tasks, Suh et al.’s [31] loss function employs the BCE loss instead of the $L1$ loss used in the Pix2Pix GAN [26]. This improved loss function of the WGAN-GP is as follows:
@@
where the penalty coefficient is $\alpha$, and the uniform sampling along a straight line between the Ground-Truth distribution $P_{y}$, and the point pairs of the generated data distribution is $P_{\hat{y}}$. $\lambda$ is used to control the relative importance of different loss terms. In the generator, the distances between the generated images and the Ground-Truth images in each color channel are minimized by the loss function $\mathbb{L}_{G}$ in Eq. (5). In the discriminator, the generated images are distinguished from the real images by the loss function $\mathbb{L}_{D}$ in Eq. (6).
@@
3.3.2 Network architecture
For the network architecture proposed in Stage-2, the four independent generators use the same encoder-decoder architecture, where the encoder extracts the features and multiple encoders of different depths use a shared decoder and subsequently concatenate the features. A skip connection is used between the encoder and the decoders. The encoder facilitates downsampling and context extraction, while the decoders undertake upsampling and amalgamate both upscaled and downsampled features. Specifically, we employ EfficientNet-B6 [36] as the encoder, which is a lightweight convolutional neural network known for its performance in image classification tasks. Instead of U-Net [21] used by Suh et al. [31], we use the improved model U-Net++ [34, 35] as the network architecture for the generator, which can further improve the model performance.
In addition, we follow the discriminator used by Suh et al. [31] as the discriminator in our network architecture due to its better generalization  [26, 29, 30, 41, 42, 43].
The detailed procedure for training the proposed GANs models is illustrated in Algorithm 1. The training process is based on our constructed training environment, which includes the optimizer selection, learning rate, and loss function settings, as described in Section 4.3. The algorithm demonstrates that the GANs models iteratively update the loss function based on the discriminator’s assessment of the trueness of the generated images. This iterative process enhances the generator’s ability to generate more realistic images.
@@ 3.4 Stage-3
In Stage-3, we employ multi-scale GANs models to generate images for both local and global binarization results to enhance the distinction between foreground and background with higher accuracy. In contrast to Stage-2, which relies on the local prediction using patch images ($224\times 224$), Stage-3 performs the global prediction on the resized original input images ($512\times 512$). Therefore, this method reduces the contextual information loss in text caused by the local predictions of Stage-2.
Figure 7 illustrates the process of two independent discriminators in Stage-3. Different from Stage-2 using 8-bit images for the local prediction, Stage-3 uses 24-bit images as the inputs of the generator. In addition, Figure 7 illustrates that the left-side input images consist of the output images from Stage-2, which are used for the local prediction. In contrast, the right-side input images represent the resized images ($512\times 512$) from the original input images for the global prediction. It is worth noting that in the binarization process of Stage-3, the network architecture of the generator remains the same as in Stage-2, differing only in the number of input channels.
@@ 4.1 Datasets
4.1.1 DIBCO Dataset
The document image binarization competition (DIBCO) provides nine datasets [44, 45, 11, 46, 12, 13, 14, 15, 16]: DIBCO 2009, H-DIBCO 2010, DIBCO 2011, H-DIBCO 2012, DIBCO 2013, H-DIBCO 2014, H-DIBCO 2016, DIBCO 2017, and H-DIBCO 2018. These datasets comprise both gray-scale and color images, including machine-printed as well as handwritten content. Participants in the competition were tasked with designing algorithms capable of extracting binarized text images from both handwritten and machine-printed images within these datasets.
@@ 4.1.3 SMADI
The synchromedia multispectral ancient document images (SMADI) dataset [1] consists of 240 multispectral images of 30 authentic historical handwritten letters. These letters, dating from the 17th to the 20th centuries, are written using iron gall ink. The original documents are deposited at the BAnQ Bibliotheque et Archives nationales du Quebec. A total of 240 images of real documents are meticulously captured, calibrated, and registered within this dataset. Each of these images is imaged with a CROMA CX MSI camera, resulting in 8 images per document.
@@ 4.1.4 Bickley Diary Dataset
The bickley diary (BD) dataset [48] has been generously donated to the Singapore Methodist Archives by Erin Bickley. This dataset comprises diaries in which the images have been affected by both light and fold damages, thereby complicating the process of identification.
@@ 4.2 Evaluation Metric
In comparison with other methods for document image enhancement and binarization, based on the case that DIBCO datasets do not provide OCR output, we use the following five metrics:
@@ 4.2.1 F-measure (FM)
where the $Recall$ is defined as $\frac{TP}{TP+FN}$; the $Precision$ is defined as $\frac{TP}{TP+FP}$, and TP, FP, and FN represent the true positive, false positive, and false negative values.
@@ 4.2.2 Pseudo-F-measure (p-FM)
where both pseudo-recall (${\rm Recall_{ps}}$) and pseudo-precision (${\rm Precision_{ps}}$) use distance weights relative to the contours of the Ground-Truth images to generate new weighted images.
@@ 4.2.3 Peak signal-to-noise ratio (PSNR)
where ${\rm V}$ denotes the difference between the foreground and background, and the mean square error (MSE) is defined as:
@@ 4.2.4 Distance reciprocal distortion (DRD)
${\rm DRD}$ represents a measure of visual distortion in the binary images, which is shown in the following equation:
@@
where ${\rm NUBN}$ corresponds to the number of uneven $8\times 8$ blocks in the Ground-Truth images, and ${\rm DRD_{k}}$ is defined as:
@@
where ${\rm DRD_{k}}$ indicates the distortion of the ${\rm k_{th}}$ flipped pixel, calculated using a $5\times 5$ normalized weight matrix ${\rm N_{w}}$. ${\rm B_{k}}$ refers to the pixel of the binary images, and ${\rm G_{k}}$ denotes the pixel of the Ground-Truth images.
@@ 4.2.5 Avg-Score (Avg)
The PSNR metric provides a brief and intuitive quantitative way to compare the difference between the generated images and the Ground-Truth images, which can be used as a supplementary metric but should not be used as the only evaluation metric. The FM and p-FM metrics are used to assess the accuracy of different methods, but higher values of these metrics do not always correspond to higher PSNR values.
Based on the results of our extensive experiments, it is common to observe that a method achieves the SOTA PSNR value, while the FM and p-FM values are lower than those of other models. To solve this problem, we follow the Avg-Score metric used by Jemni et al. [49] to compare the combined performance of different methods, as shown in the following equation:
@@
Table 1: Quantitative comparison of PSNR values (dB) of images processed by different techniques with the corresponding Ground-Truth images on different training sets. The highest and 2nd highest PSNR values are in red and blue colors, respectively.
@@
Original 65.9812 62.7183 63.6814 59.3267 65.0543 62.5198 63.2136
LL (DWT) 56.8869 62.9102 57.4896 50.4236 51.6242 57.1778 56.0854
HL (DWT) 49.8898 49.7108 49.9529 50.8580 50.0277 49.7025 50.0236
LH (DWT) 49.9061 49.7136 49.9578 50.8483 50.0311 49.6946 50.0253
HH (DWT) 49.0228 48.8979 49.2434 50.0180 49.2218 48.9594 49.2272
LL (DWT&Norm) 70.2888 77.0135 74.3312 65.1956 71.4992 69.4219 71.2917
HL (DWT&Norm) 56.8078 58.2430 59.3027 56.6744 57.1121 56.7830 57.6497
LH (DWT&Norm) 56.0911 58.7330 58.9662 57.0237 57.9381 56.9083 57.6101
HH (DWT&Norm) 63.5701 64.5478 63.4654 61.8377 61.6075 61.7661 62.7991
Table 2: Quantitative comparison (FM/p-FM/PSNR/DRD/Avg) of the proposed methods with different patch sizes ($224\times 224$ vs. $256\times 256$) on multiple benchmark datasets.
DIBCO 2011 DIBCO 2013 H-DIBCO 2014 H-DIBCO 2016 DIBCO 2017 H-DIBCO 2018 Mean Values
@@
Figure 8: Visualization examples of the results of the proposed method at each stage: (a) the original input images, (b) the LL subband images processed by Haar wavelet transform, (c) the LL subband images processed by Haar wavelet transform with normalization (Stage-1), (d) the enhanced images produced using our image enhancement method (Stage-2), (e) the binarized images obtained by combining local and global predictions (Stage-3), (f) the Ground-Truth images.
Table 3: Quantitative comparison of performance evaluation (FM/p-FM/PSNR/DRD/Avg) of the proposed method and other methods for document image binarization on multiple benchmark datasets. Best and 2nd best performance are in red and blue colors, respectively. The data for the experimental results of other methods are obtained from [31, 50].
@@
(f) H-DIBCO 2018
Figure 9: Example of document image binarization results from the input image “HW1” of DIBCO 2011 dataset by different methods: (a) original input image, (b) Otsu’s [6], (c) Niblack’s [5], (d) Sauvola et al.’s [7], (e) Vo et al.’s [23], (f) He et al.’s [24], (g) Suh et al.’s [31], (h) Ours, and (i) Ground-Truth image.
Figure 10: Example of document image binarization results from input image “HW5” of DIBCO 2013 dataset by different methods: (a) original input image, (b) Otsu’s [6], (c) Niblack’s [5], (d) Sauvola et al.’s [7], (e) Vo et al.’s [23], (f) He et al.’s [24], (g) Suh et al.’s [31], (h) Ours, and (i) Ground-Truth image.
@@ 4.3.1 Data preparation
The model performance of the neural network depends on the quality and quantity of its training data. In general, the more training data improves the better model performance. To fairly compare the performance of the proposed method with other methods, we follow the strategy described in [23, 24, 31] to construct the training and test sets. Specifically, the training set includes a total of 143 images from DIBCO 2009 [44], H-DIBCO 2010 [45], H-DIBCO 2012 [46], PHIBD [47], SMADI [1], and BD [48] dataset. While the test sets comprise a total of 82 images from DIBCO 2011 [11], DIBCO 2013 [12], H-DIBCO 2014 [13], H-DIBCO 2016 [14], DIBCO 2017 [15], and H-DIBCO 2018 [16].
To ensure a fair comparison of the performance of all methods, for the proposed method and other methods, we apply the same data augmentation techniques to the training data. In the training process of the GANs models of Stage-2, we split the original input images into patches of $224\times 224$ size, and these patches are sampled with scale factors of ${0.75,1,1.25,1.5}$. In addition, we rotate the image by $270\degree$ to obtain more training data. In contrast, for training GAN model on resized $512\times 512$ images of Stage-3, we employ scaling without rotation for data augmentation. To extend the training data, we employ the rotation augmentation technique during the training process of this model, including $90\degree$, $180\degree$ and $270\degree$ rotations, horizontal and vertical flips.
@@ 4.3.2 Training
The experiments are conducted using the PyTorch framework in Python. We use single NVIDIA RTX 4090 GPU for all the experiments.
In this work, we employ the generator consisting of EfficientNet-B6 [36] with U-Net++ [34, 35]. About the training hyper-parameters, the settings of Stage 2 and Stage 3 are similar, with the only difference in the number of training epochs. Specifically, for the network of the global prediction we set it to 150 epochs and for the remaining networks we set it to 10 epochs. We select the Adam [51] optimizer, with an initial learning rate set to $2\times 10^{-4}$. In addition, we set $\beta_{1}=0.5$ for the generator, and $\beta_{2}=0.999$ for the discriminator.
@@ 4.3.3 Pre-training
Due to the limited data available, to improve the training efficiency, we choose EfficientNet-B6 [36] as the encoder of the generator, utilizing the pre-trained weights on the ImageNet dataset [52].
@@ 4.4 Comparison with other Methods
In our previous conference paper [53], we compared different deep learning-based methods and found that some of these methods [10, 28, 29] employ “leave-one-out” strategy for constructing the training set. For instance, when the DIBCO 2013 [12] dataset is used as the test set, all other DIBCO datasets are used as the training set. Although our method still shows excellent performance in such case of [53], we consider it is unfair to compare them directly because our model uses a much smaller training set. Therefore, we exclude these methods from the methods we compare in this paper, and retain only those methods [5, 6, 7, 17, 18, 23, 24, 31] that use the same training set for comparison.
@@ 4.5 Image Enhancement Experiment
The previous conference paper [53] of this work presents ablation study on the DWT with normalization. Based on this, it is better to intuitively show that, compared to the binarization results of the original input images, the binarization results of the images processed by the DWT with normalization are closer to the Ground-Truth images.
While the mathematical theory [54] suggests that the images processed by the Haar wavelet transform can retain contour information and reduce noise more efficiently, it is important to demonstrate their impact on experimental results. In Stage-1 of this work, we set the threshold to 0.5 to binarize different single-channel $224\times 224$ patch images (red, green, and blue) to obtain the corresponding binarization results. In the experiment of this section, we apply the above image binarization to the original input images, the images only processed by the DWT, and the images processed by the DWT with normalization. Subsequently, we calculate the PSNR values of the binarized images by different techniques compared to the Ground-Truth images, and the results are presented in Table 1. Compared with the PSNR mean value of the original input images of 63.2136dB, the PSNR mean values of the four subband images (LL, HL, LH, HH) only processed by the DWT are all lower, which are 56.0854dB, 50.0236dB, 50.0253dB, and 49.2272dB, respectively. In addition, the PSNR mean values of the three subband images (HL, LH, LL) processed by the DWT with normalization are lower than that of the original input images, which are 57.6497dB, 57.6101dB, and 62.7991dB, respectively, while the PSNR mean value of the LL subband images processed by the DWT with normalization reaches 71.2917dB, which means they are closer to the Ground-Truth images. The results of this experiment show that the choice of the DWT with normalization is more reasonable than only the DWT for processing the images.